# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 4 Pre- and Post-conditions Detection
In pre- and post-conditions detection, it is necessary to identify the initial and termination conditions. In pre-conditions detection, recognize objects and observe the affordances of objects. In post-conditions detection, identify whether a task has been successfully executed and provide reasons for task failure after skill execution. Currently, there are few papers focusing on identifying termination conditions. Cui et al. (2022) utilizes CLIP to compare the target’s text or image with the termination environment to determine the success of task execution. Few articles are found in this study that address the output of task failure reasons after skill execution. RobotGPT (Jin et al. (2024)) analysis task failure utilizes the positions of manipulated objects after execution, but task failure should be determined during execution. AHA (Duan et al. (2024)) uses a large number of robotic failure trajectories to fine-tune the VLM. The fine-tuned VLM leverages keyframe trajectory images and task descriptions from the robot’s current task execution process to detect failures and provide detailed, adaptable failure explanations. Therefore, this section focuses on literature discussing foundation models in pre-conditions detection including object affordance and object recognition.
@@ 4.1 Object Affordance
The affordances associated with an object represent the range of manipulations that the object affords the robot (Gibson (2014)). Early approaches addressed the issue by treating it as a supervised task (Kokic et al. (2017)). However, the process of annotating datasets is laborious and time-consuming, making it impractical to exhaustively cover all geometric information present in real-world environments. Consequently, researchers are exploring the application of reinforcement learning, enabling robots to collect data and train affordance perception modules through continuous exploration (Wu et al. (2021)). Nevertheless, current reinforcement learning methods are trained in simulated environments, leading to a significant sim-to-real gap. To address these challenges, researchers propose training the affordance perception module using videos of human interactions in real-world scenarios (Ye et al. (2023b); Bahl et al. (2023)).
@@
For supervised learning methods, GraspGPT (Tang et al. (2023)) utilizes LLM outputs for object class descriptions and task descriptions. Object class descriptions detail the geometric shapes of each part of an object, while task descriptions outline the desired affordances for task execution, such as the types of manipulation actions to be taken. Integrating both components into the task-oriented grasp evaluator enhances the quality of the generated grasp pose. 3DAP (Nguyen et al. (2023)) utilizes the text encoder of LLM for feature extraction. The extracted features from desired affordances text are inputted into both the affordance detection module and pose generation module. This enhances the quality of the predicted affordance map and the generated pose.
@@
In reinforcement learning, ATLA (Ren et al. (2023b)) utilizes GPT-3 to generate language descriptions of tools. These descriptions are then inputted into a pre-trained BERT model to obtain representations. The extracted features are finally fed into the SAC network module. Meta-learning techniques are employed to enhance the learning efficiency for the use of new tools. Xu et al. (2023a) employ CLIP’s text and image encoders to extract features from language instruction and scene image, improving the quality of grasp pose generation in the SAC module.
@@
The methods mentioned above utilize foundation models to assist other learning methods in improving affordance maps or grasp poses. There are also direct approaches using foundation models to generate affordance maps and grasp poses. PartSLIP (Liu et al. (2023d)) converts 3D point clouds into 2D rendering images and inputs multi-view 2D images and textual descriptions of object parts into GLIP for object parts detection, ultimately fusing 2D bounding boxes into 3D segmentation to generate affordance maps. However, PartSLIP requires manual definition text prompts and additional algorithms to convert 2D boxes back to 3D regions. UAD (Tang et al. (2025)) clusters object points into fine-grained semantic regions based on pixel-wise features extracted from multi-view rendered images using DINOv2. It then queries the VLM to generate a set of task instructions and associates these instructions with the most relevant clustered region to construct the affordance map. LAN-grasp (Mirjalili et al. (2023)) inputs human instruction into LLM, utilizing its prior knowledge to output the shape of part to be grasped. These shapes, along with the object’s 2D image, are then inputted into VLM to detect the bounding box for the grasping part. Finally, the bounding box and the point cloud from object 3D reconstruction are inputted into the grasp planner to generate grasp poses.
@@
Figure 4: Foundation Models for Pre-conditions Detection. As for object affordance, the main approaches of task-oriented grasp are supervised learning and reinforcement learning. Both methods utilize LLM to generate object part-level description and desired affordance description in task instruction, then fuse tokens and features into the original network through language encoder and image encoder to output task-oriented grasp pose (Tang et al. (2023); Ren et al. (2023b)). In reinforcement learning, it is possible to choose between a LLM language encoder with a custom-designed image encoder, or a VLM language encoder with a VLM image encoder. When selecting the LLM language encoder with a custom image encoder, the LLM language encoder should be frozen, and the custom image encoder should be trained (Ren et al. (2023b)). When using the VLM language encoder with the VLM image encoder, both encoders are typically frozen (Xu et al. (2023a)). Direct using foundation method utilizes LLM to generate object part-level description and desired affordance description according to task instruction. VLM marks out the part of the object to grasp in the image based on the description (Liu et al. (2023d)). As for object recognition, the representation learning methods in state perception mainly include contrastive learning (Radford et al. (2021)), distillation-based learning (Caron et al. (2021)), and masked autoencoder learning (Radosavovic et al. (2023)). Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse, the fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). Multimodal representation learning focuses primarily on multimodal alignment (Xue et al. (2023b); Tatiya et al. (2023)). Training the encoder with large-scale data and parameters has facilitated open-set perception, including tasks such as open-set detection, open-set segmentation. For instance, SAM (Kirillov et al. (2023)) utilizes the MAE (He et al. (2022)), ViLD (Gu et al. (2021)) employs the CLIP (Radford et al. (2021)).
@@ 4.2 Object Recognition
Object recognition can be categorized into two types: passive perception and active perception. Compared to passive perception, active perception adjusts the perspective to the areas of interest (Kroemer et al. (2021)). Then, modeling manipulation tasks and generalizing manipulation skills necessitate representations of both the robot’s environment and the manipulated objects. These representations form the foundation for skill hierarchies, pre- and post-condition detection, skill learning, and transition model learning.
@@
The Vision Transformers (ViTs) and similar attention-based neural networks have recently achieved state-of-the-art performance on numerous computer vision benchmarks (Han et al. (2022); Khan et al. (2022); Zhai et al. (2022)) and the scaling of ViTs has driven breakthrough capability for vision models (Dehghani et al. (2023)). The development of visual backbones not only advances pre-trained visual representations but also accelerates the progress of open-set perception tasks, such as segmentation and detection.
@@
As for pre-trained visual representations, the algorithms mentioned have various training objectives: for instance, contrastive methods like Vi-PRoM (Caron et al. (2021)), R3M (Nair et al. (2022)), VIP (Ma et al. (2022)), CLIP (Radford et al. (2021)), LIV (Ma et al. (2023a)); distillation-based methods such as DINO (Caron et al. (2021)); or masked autoencoder methods like MAP (Radosavovic et al. (2023)), MAE (He et al. (2022)). The primary datasets utilized comprise the CLIP dataset (Radford et al. (2021)), consisting of 400 million (image, text) pairs sourced from the internet, along with ImageNet (Deng et al. (2009)), Ego4D (Grauman et al. (2022)), and EgoNet (Jing et al. (2023)).
@@
Pre-trained visual representations have high transfer ability to policy learning (Xiao et al. (2022b); Yang et al. (2023c)), but visual representation involves not just recognizing spatial features but also understanding semantic features. Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse (Karamcheti et al. (2023)). The fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). The loss function achieves a balanced trade-off between these two aspects. To compare different pre-trained visual representations, benchmarks are established by CORTEXBENCH (Majumdar et al. (2023)) and EmbCLIP (Khandelwal et al. (2022)) to assess which model could provide a better ”artificial visual cortex” for manipulation tasks. However, the models included in these benchmarks are still not comprehensive enough.
@@
The aforementioned pre-trained visual representations mainly involve the extraction of features from 2D images. The experience of learning representations on 2D images can also be extended to other modalities. For the object point cloud modality, ULIP (Xue et al. (2023a)) and ULIP2 (Xue et al. (2023b)) employ contrastive learning to align features between point clouds and text-images. Point-BERT (Yu et al. (2022)) uses the masked autoencoding method to learn point cloud features by reconstructing point clouds. GeDi (Poiesi and Boscaini (2022)) uses a contrastive learning approach to extract general and distinctive 3D local geometric information. In the haptic modality, MOSAIC (Tatiya et al. (2023)) utilizes contrastive learning to train the haptic encoder.
@@
As for segmentation, SAM (Kirillov et al. (2023)) develops a transformer-based architecture and creates the largest segmentation dataset, with over 1 billion masks from 11 million images. The model is adaptable and enables zero-shot transfer to new tasks and image distributions. Fast-SAM (Zhao et al. (2023b)) and Faster-SAM (Zhang et al. (2023a)) aim to improve the training and inference speed of the network by enhancing its network structure. TAM (Yang et al. (2023a)) merges SAM (Yang et al. (2023a)) and XMem (Cheng and Schwing (2022)) for high-performance interactive tracking and segmentation in videos.
@@
As for detection, traditional detection models are usually confined to a narrow range of semantic categories because of the cost and time involved in gathering localized training data within extensive or open-label domains. However, advancements in language encoders and contrastive image-text training enable open-set detection. Researchers integrate language into a closed-set detector to generalize open-set concepts, detecting various classes through language generalization despite being trained solely on existing bounding box annotations, such as OWL-ViT (Minderer et al. (2022)), Grounding-DINO (Liu et al. (2023f)), OVD (Zareian et al. (2021)), ViLD (Gu et al. (2021)), DetCLIP (Yao et al. (2022a)).
@@
Deploying such models in open-set detection presents a significant challenge, primarily because even slight alterations in prompting can greatly impact performance. Fine-tuning can enhance a foundation model’s understanding of prompting. However, foundation models are often over-parameterized, leading to slow training processes. COOP (Zhou et al. (2022)) maps prompting to a set of learnable vectors, which can be optimized through network training. In CLIP-Adapter (Gao et al. (2024)), two extra linear layers are appended after the final layer of either the vision or language backbone to enable efficient few-shot transfer learning through fine-tuning.
@@
The method for open-set detection on 2D images can be extended to the research direction of open-set detection on 3D point clouds. PointCLIP (Zhang et al. (2022b)) utilizes pre-trained CLIP to extract multi-view depth image features of point cloud, then compares the extracted features with textual features to identify the point cloud category.
@@ 5 Hierarchy of skills
The skill hierarchy is closely related to the field of task and motion planning (TAMP). TAMP aims to address high-level instructions by organizing tasks in a sequence that ensures dynamic feasibility (Guo et al. (2023)). There are three main types of classical TAMP methods: constraint-based TAMP, sampling-based TAMP, and optimization-based TAMP (Zhao et al. (2024)). Constraint-based and sampling-based TAMP define the problem with goal conditions. Unlike optimization-based TAMP, these approaches often cannot assess or compare the quality of the generated plan or final state due to the lack of objective functions, such as when the goal is to pour as much water as possible into the cup (Zhao et al. (2024); Zhang et al. (2022a)). However, optimization-based TAMP is sensitive to the initial conditions and goal setup of the problem (Zhao et al. (2024)).
@@
The scalability of classical TAMP methods is often constrained by the tree search problem size for complex tasks and the computational cost of evaluating heuristics and optimal trajectories (Zhao et al. (2024)). Integrating learning-based approaches into TAMP enables informed decision-making based on prior examples and experiences and improves flexibility and generalizability (Guo et al. (2023)). Models for skill hierarchy can be trained using text or videos, similar to how humans learn assembly procedures from instructional manuals or tutorial videos. As for tutorial videos, VLaMP (Patel et al. (2023) and SeeDo (Wang et al. (2024a)) use trained models to understand human video operations and HourVideo (Chandrasegaran et al. (2024)) proposes a benchmark dataset specifically designed for hour-long video-language understanding.
@@
Traditional TAMP’s domain representations are usually manually specified by expert users such as PDDL (Silver et al. (2022)). However, LLMs have been explored for processing and interpreting natural language inputs (Huang et al. (2022)). They offer a novel approach to encoding the planning domain in a more intuitive and accessible way. Furthermore, LLMs’ acquisition of world knowledge and commonsense reasoning has the potential to improve the scalability and generalizability of skill hierarchy tasks (Vemprala et al. (2023); Jansen (2020); Driess et al. (2023)). Various benchmarks such as PlanBench (Valmeekam et al. (2023)) can assess the planning and reasoning capability of LLMs.
@@
LLMs possess a notable limitation: they lack practical experience, hindering their utility for decision-making within a specific context, so the output of LLMs often cannot be translated into executable actions for the robot. Huang et al. (2022) first use pre-trained causal LLM to break down high-level tasks into logical mid-level action plans. Then, a pre-trained masked LLM is employed to convert mid-level action plans into admissible actions. However, prompts usually require the context of the robot’s capability, its current state, and the environment. At the same time, LLMs are considered ‘forgetful’ and don’t treat information in the system prompt as absolute. Despite efforts to reinforce task constraints in the objective prompt and extract numerical task contexts from the system prompt, storing them in data structures, errors caused by LLM forgetfulness remain unresolved (Chen and Huang (2023)).
@@
To address the aforementioned issues, SayCan (Ahn et al. (2022)) scores pre-trained tasks based on prompting and observation images, generating the task sequence with the highest score. Saycan provides a paradigm for generating action sequences using LLM, but there are still some drawbacks: 1) The generated action sequences do not incorporate user preferences. 2) Safety regulations are not adequately addressed. 3) The limitation of the skill library. 4) LLM focuses solely on reasoning when generating action sequences, neglecting feedback on action execution. 5) The limitation of scene grounding. GD (Huang et al. (2023d)) proposes a paradigm to address the aforementioned issues by not only scoring the generated action sequence using LLM but also introducing a grounded function model for scoring the generated action sequence. The grounded function model encompasses token-conditioned robotic functions, such as affordance functions that capture the abilities of a robot based on its embodiment, safety functions, and more. This approach tackles drawbacks by designing grounded functions, avoiding fine-tuning in LLM.
@@
Regarding user preferences, TidyBot (Wu et al. (2023b)) trains LLM by collecting users’ preference data, enabling the trained LLM to choose behaviors that better align with user preferences. As for safety regulations, Yang et al. (2023d) incorporate ISO 61508, a global standard for safely deploying robots in industrial factory settings, into the constraints of the action sequence generation. As for the skill library, BOSS (Zhang et al. (2023c)) suggests using LLMs’ rich knowledge to guide skills chaining in the skill library, aiming to create new skills through combinations. RoboGen (Wang et al. (2023e)) employs generative models to create new skill task scenarios, then utilizes either reinforcement learning or gradient optimization methods to automatically learn new skills based on the reward function generated by the LLM. As for action execution feedback, REACT (Yao et al. (2022b)), COWP (Ding et al. (2022)), LLM-Planner (Song et al. (2023)), CoPAL (Joublin et al. (2023)) and PROGPROMPT (Singh et al. (2023)) provide feedback on robot action execution to LLMs. This allows LLMs to adjust action sequences based on execution status, creating a closed-loop process for generating action sequences.
@@
Figure 5: Foundation Models for Hierarchy of Skills. 1) Utilize human operation video to learn the skill sequence for task execution, decompose the video of the user’s progress so far into observations and human actions through segmentation, and input them along with task instruction into a pre-trained language model to predict the next step (Patel et al. (2023)). 2) LLM scores the skills in the skill library based on task instruction and the skills already executed, and the value function also scores the skills in the skill library based on observation images. The highest-scoring skill, obtained by multiplying the two scores, is selected as the next step (Ahn et al. (2022)). The value function can consider multiple factors such as affordance, safety, user preference, and more (Huang et al. (2023d)), and these considerations can also be fine-tuning LLM (Wu et al. (2023b)). 3) LLM assists the classical planner by translating task instruction into PDDL descriptions, sending them to the classical planner to generate a PDDL plan, and then translating the PDDL plan into a natural language plan using LLM (Liu et al. (2023a)).
@@
As for the limitation of scene grounding, LLMs need to inquire about the scene representation to determine the availability, relationship and location of objects. NLMap (Chen et al. (2023a)) proposes an open-vocabulary, queryable semantic representation map built on ViLD and CLIP. This map outputs the pose of related objects based on task instruction, which are then handed over to the LLM for planning. Text2Motion (Lin et al. (2023)) incorporates a geometric value function on top of the value function, enabling the robot to select actions that adhere to geometric constraints based on scene descriptions. Xu et al. (2023b) explore the possibility of teaching robots to creatively utilize tools within scenarios, which involve implicit physical limitations and require long-term planning. VILA (Hu et al. (2023a)) seamlessly incorporates perceptual data into ChatGPT-4V for its reasoning and planning processes, facilitating a deep comprehension of common sense knowledge within the visual domain, encompassing spatial arrangements and object characteristics. PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to enhance its understanding of physical object attributes, such as material. This integration of a physically informed VLM into an interactive framework with a LLM enhances task planning performance in tasks incorporating instruction related to physical object attributes. SpatialVLM (Chen et al. (2024b)) and 3D-LLM (Hong et al. (2023a)) utilizes a 2D pre-trained VLM to train on collected 3D datasets, enhancing capabilities related to 3D tasks while maintaining the abilities of previous tasks.
@@
The hierarchy of skills possessed by LLMs or VLMs can be applied not only to single agent but also to multiple agents. SMART-LLM (Kannan et al. (2023)) utilizes LLM for the hierarchy of skills and allocates each task to every agent through the task assignment module.

### v7

@@ 4 Pre- and Post-conditions Detection
In pre- and post-conditions detection, it is necessary to identify the initial and termination conditions. In pre-conditions detection, recognize objects and observe the affordances of objects. In post-conditions detection, identify whether a task has been successfully executed and provide reasons for task failure after skill execution. Currently, there are few papers focusing on identifying termination conditions. cui2022can utilizes CLIP to compare the target’s text or image with the termination environment to determine the success of task execution. Few articles are found in this study that address the output of task failure reasons after skill execution. RobotGPT (jin2024robotgpt) analysis task failure utilizes the positions of manipulated objects after execution, but task failure should be determined during execution. AHA (duan2024aha) uses a large number of robotic failure trajectories to fine-tune the VLM. The fine-tuned VLM leverages keyframe trajectory images and task descriptions from the robot’s current task execution process to detect failures and provide detailed, adaptable failure explanations. Therefore, this section focuses on literature discussing foundation models in pre-conditions detection including object affordance and object recognition.
@@ 4.1 Object Affordance
The affordances associated with an object represent the range of manipulations that the object affords the robot (gibson2014ecological). Early approaches addressed the issue by treating it as a supervised task (kokic2017affordance). However, the process of annotating datasets is laborious and time-consuming, making it impractical to exhaustively cover all geometric information present in real-world environments. Consequently, researchers are exploring the application of reinforcement learning, enabling robots to collect data and train affordance perception modules through continuous exploration (wu2021vat). Nevertheless, current reinforcement learning methods are trained in simulated environments, leading to a significant sim-to-real gap. To address these challenges, researchers propose training the affordance perception module using videos of human interactions in real-world scenarios (ye2023affordance; bahl2023affordances).
@@
For supervised learning methods, GraspGPT (tang2023graspgpt) utilizes LLM outputs for object class descriptions and task descriptions. Object class descriptions detail the geometric shapes of each part of an object, while task descriptions outline the desired affordances for task execution, such as the types of manipulation actions to be taken. Integrating both components into the task-oriented grasp evaluator enhances the quality of the generated grasp pose. 3DAP (nguyen2023language) utilizes the text encoder of LLM for feature extraction. The extracted features from desired affordances text are inputted into both the affordance detection module and pose generation module. This enhances the quality of the predicted affordance map and the generated pose.
@@
In reinforcement learning, ATLA (ren2023leveraging) utilizes GPT-3 to generate language descriptions of tools. These descriptions are then inputted into a pre-trained BERT model to obtain representations. The extracted features are finally fed into the SAC network module. Meta-learning techniques are employed to enhance the learning efficiency for the use of new tools. xu2023joint employ CLIP’s text and image encoders to extract features from language instruction and scene image, improving the quality of grasp pose generation in the SAC module.
@@
The methods mentioned above utilize foundation models to assist other learning methods in improving affordance maps or grasp poses. There are also direct approaches using foundation models to generate affordance maps and grasp poses. PartSLIP (liu2023partslip) converts 3D point clouds into 2D rendering images and inputs multi-view 2D images and textual descriptions of object parts into GLIP for object parts detection, ultimately fusing 2D bounding boxes into 3D segmentation to generate affordance maps. However, PartSLIP requires manual definition text prompts and additional algorithms to convert 2D boxes back to 3D regions. UAD (tang2025uad) clusters object points into fine-grained semantic regions based on pixel-wise features extracted from multi-view rendered images using DINOv2. It then queries the VLM to generate a set of task instructions and associates these instructions with the most relevant clustered region to construct the affordance map. LAN-grasp (mirjalili2023lan) inputs human instruction into LLM, utilizing its prior knowledge to output the shape of part to be grasped. These shapes, along with the object’s 2D image, are then inputted into VLM to detect the bounding box for the grasping part. Finally, the bounding box and the point cloud from object 3D reconstruction are inputted into the grasp planner to generate grasp poses.
@@
Figure 4: Foundation Models for Pre-conditions Detection. As for object affordance, the main approaches of task-oriented grasp are supervised learning and reinforcement learning. Both methods utilize LLM to generate object part-level description and desired affordance description in task instruction, then fuse tokens and features into the original network through language encoder and image encoder to output task-oriented grasp pose (tang2023graspgpt; ren2023leveraging). In reinforcement learning, it is possible to choose between a LLM language encoder with a custom-designed image encoder, or a VLM language encoder with a VLM image encoder. When selecting the LLM language encoder with a custom image encoder, the LLM language encoder should be frozen, and the custom image encoder should be trained (ren2023leveraging). When using the VLM language encoder with the VLM image encoder, both encoders are typically frozen (xu2023joint). Direct using foundation method utilizes LLM to generate object part-level description and desired affordance description according to task instruction. VLM marks out the part of the object to grasp in the image based on the description (liu2023partslip). As for object recognition, the representation learning methods in state perception mainly include contrastive learning (radford2021learning), distillation-based learning (caron2021emerging), and masked autoencoder learning (radosavovic2023real). Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse, the fusion of masked autoencoder and contrastive learning is employed in both Voltron (karamcheti2023language) and iBOT (zhou2021ibot). Multimodal representation learning focuses primarily on multimodal alignment (xue2023ulip2; tatiya2023mosaic). Training the encoder with large-scale data and parameters has facilitated open-set perception, including tasks such as open-set detection, open-set segmentation. For instance, SAM (kirillov2023segment) utilizes the MAE (he2022masked), ViLD (gu2021open) employs the CLIP (radford2021learning).
@@ 4.2 Object Recognition
Object recognition can be categorized into two types: passive perception and active perception. Compared to passive perception, active perception adjusts the perspective to the areas of interest (kroemer2021review). Then, modeling manipulation tasks and generalizing manipulation skills necessitate representations of both the robot’s environment and the manipulated objects. These representations form the foundation for skill hierarchies, pre- and post-condition detection, skill learning, and transition model learning.
@@
The Vision Transformers (ViTs) and similar attention-based neural networks have recently achieved state-of-the-art performance on numerous computer vision benchmarks (han2022survey; khan2022transformers; zhai2022scaling) and the scaling of ViTs has driven breakthrough capability for vision models (dehghani2023scaling). The development of visual backbones not only advances pre-trained visual representations but also accelerates the progress of open-set perception tasks, such as segmentation and detection.
@@
As for pre-trained visual representations, the algorithms mentioned have various training objectives: for instance, contrastive methods like Vi-PRoM (caron2021emerging), R3M (nair2022r3m), VIP (ma2022vip), CLIP (radford2021learning), LIV (ma2023liv); distillation-based methods such as DINO (caron2021emerging); or masked autoencoder methods like MAP (radosavovic2023real), MAE (he2022masked). The primary datasets utilized comprise the CLIP dataset (radford2021learning), consisting of 400 million (image, text) pairs sourced from the internet, along with ImageNet (deng2009imagenet), Ego4D (grauman2022ego4d), and EgoNet (jing2023exploring).
@@
Pre-trained visual representations have high transfer ability to policy learning (xiao2022masked; yang2023moma), but visual representation involves not just recognizing spatial features but also understanding semantic features. Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse (karamcheti2023language). The fusion of masked autoencoder and contrastive learning is employed in both Voltron (karamcheti2023language) and iBOT (zhou2021ibot). The loss function achieves a balanced trade-off between these two aspects. To compare different pre-trained visual representations, benchmarks are established by CORTEXBENCH (majumdar2023we) and EmbCLIP (khandelwal2022simple) to assess which model could provide a better ”artificial visual cortex” for manipulation tasks. However, the models included in these benchmarks are still not comprehensive enough.
@@
The aforementioned pre-trained visual representations mainly involve the extraction of features from 2D images. The experience of learning representations on 2D images can also be extended to other modalities. For the object point cloud modality, ULIP (xue2023ulip) and ULIP2 (xue2023ulip2) employ contrastive learning to align features between point clouds and text-images. Point-BERT (yu2022point) uses the masked autoencoding method to learn point cloud features by reconstructing point clouds. GeDi (poiesi2022learning) uses a contrastive learning approach to extract general and distinctive 3D local geometric information. In the haptic modality, MOSAIC (tatiya2023mosaic) utilizes contrastive learning to train the haptic encoder.
@@
As for segmentation, SAM (kirillov2023segment) develops a transformer-based architecture and creates the largest segmentation dataset, with over 1 billion masks from 11 million images. The model is adaptable and enables zero-shot transfer to new tasks and image distributions. Fast-SAM (zhao2023fast) and Faster-SAM (zhang2023faster) aim to improve the training and inference speed of the network by enhancing its network structure. TAM (yang2023track) merges SAM (yang2023track) and XMem (cheng2022xmem) for high-performance interactive tracking and segmentation in videos.
@@
As for detection, traditional detection models are usually confined to a narrow range of semantic categories because of the cost and time involved in gathering localized training data within extensive or open-label domains. However, advancements in language encoders and contrastive image-text training enable open-set detection. Researchers integrate language into a closed-set detector to generalize open-set concepts, detecting various classes through language generalization despite being trained solely on existing bounding box annotations, such as OWL-ViT (minderer2022simple), Grounding-DINO (liu2023grounding), OVD (zareian2021open), ViLD (gu2021open), DetCLIP (yao2022detclip).
@@
Deploying such models in open-set detection presents a significant challenge, primarily because even slight alterations in prompting can greatly impact performance. Fine-tuning can enhance a foundation model’s understanding of prompting. However, foundation models are often over-parameterized, leading to slow training processes. COOP (zhou2022learning) maps prompting to a set of learnable vectors, which can be optimized through network training. In CLIP-Adapter (gao2024clip), two extra linear layers are appended after the final layer of either the vision or language backbone to enable efficient few-shot transfer learning through fine-tuning.
@@
The method for open-set detection on 2D images can be extended to the research direction of open-set detection on 3D point clouds. PointCLIP (zhang2022pointclip) utilizes pre-trained CLIP to extract multi-view depth image features of point cloud, then compares the extracted features with textual features to identify the point cloud category.
@@ 5 Hierarchy of skills
The skill hierarchy is closely related to the field of task and motion planning (TAMP). TAMP aims to address high-level instructions by organizing tasks in a sequence that ensures dynamic feasibility (guo2023recent). There are three main types of classical TAMP methods: constraint-based TAMP, sampling-based TAMP, and optimization-based TAMP (zhao2024survey). Constraint-based and sampling-based TAMP define the problem with goal conditions. Unlike optimization-based TAMP, these approaches often cannot assess or compare the quality of the generated plan or final state due to the lack of objective functions, such as when the goal is to pour as much water as possible into the cup (zhao2024survey; zhang2022task). However, optimization-based TAMP is sensitive to the initial conditions and goal setup of the problem (zhao2024survey).
@@
The scalability of classical TAMP methods is often constrained by the tree search problem size for complex tasks and the computational cost of evaluating heuristics and optimal trajectories (zhao2024survey). Integrating learning-based approaches into TAMP enables informed decision-making based on prior examples and experiences and improves flexibility and generalizability (guo2023recent). Models for skill hierarchy can be trained using text or videos, similar to how humans learn assembly procedures from instructional manuals or tutorial videos. As for tutorial videos, VLaMP (patel2023pretrained and SeeDo (wang2024vlm) use trained models to understand human video operations and HourVideo (chandrasegaran2024hourvideo) proposes a benchmark dataset specifically designed for hour-long video-language understanding.
@@
Traditional TAMP’s domain representations are usually manually specified by expert users such as PDDL (silver2022pddl). However, LLMs have been explored for processing and interpreting natural language inputs (huang2022language). They offer a novel approach to encoding the planning domain in a more intuitive and accessible way. Furthermore, LLMs’ acquisition of world knowledge and commonsense reasoning has the potential to improve the scalability and generalizability of skill hierarchy tasks (vemprala2023chatgpt; jansen2020visually; palme). Various benchmarks such as PlanBench (valmeekam2023planbench) can assess the planning and reasoning capability of LLMs.
@@
LLMs possess a notable limitation: they lack practical experience, hindering their utility for decision-making within a specific context, so the output of LLMs often cannot be translated into executable actions for the robot. huang2022language first use pre-trained causal LLM to break down high-level tasks into logical mid-level action plans. Then, a pre-trained masked LLM is employed to convert mid-level action plans into admissible actions. However, prompts usually require the context of the robot’s capability, its current state, and the environment. At the same time, LLMs are considered ‘forgetful’ and don’t treat information in the system prompt as absolute. Despite efforts to reinforce task constraints in the objective prompt and extract numerical task contexts from the system prompt, storing them in data structures, errors caused by LLM forgetfulness remain unresolved (chen2023forgetful).
@@
To address the aforementioned issues, SayCan (ahn2022can) scores pre-trained tasks based on prompting and observation images, generating the task sequence with the highest score. Saycan provides a paradigm for generating action sequences using LLM, but there are still some drawbacks: 1) The generated action sequences do not incorporate user preferences. 2) Safety regulations are not adequately addressed. 3) The limitation of the skill library. 4) LLM focuses solely on reasoning when generating action sequences, neglecting feedback on action execution. 5) The limitation of scene grounding. GD (huang2023grounded) proposes a paradigm to address the aforementioned issues by not only scoring the generated action sequence using LLM but also introducing a grounded function model for scoring the generated action sequence. The grounded function model encompasses token-conditioned robotic functions, such as affordance functions that capture the abilities of a robot based on its embodiment, safety functions, and more. This approach tackles drawbacks by designing grounded functions, avoiding fine-tuning in LLM.
@@
Regarding user preferences, TidyBot (wu2023tidybot) trains LLM by collecting users’ preference data, enabling the trained LLM to choose behaviors that better align with user preferences. As for safety regulations, yang2023plug incorporate ISO 61508, a global standard for safely deploying robots in industrial factory settings, into the constraints of the action sequence generation. As for the skill library, BOSS (zhang2023bootstrap) suggests using LLMs’ rich knowledge to guide skills chaining in the skill library, aiming to create new skills through combinations. RoboGen (wang2023robogen) employs generative models to create new skill task scenarios, then utilizes either reinforcement learning or gradient optimization methods to automatically learn new skills based on the reward function generated by the LLM. As for action execution feedback, REACT (yao2022react), COWP (ding2022robot), LLM-Planner (song2023llm), CoPAL (joublin2023copal) and PROGPROMPT (singh2023progprompt) provide feedback on robot action execution to LLMs. This allows LLMs to adjust action sequences based on execution status, creating a closed-loop process for generating action sequences.
@@
Figure 5: Foundation Models for Hierarchy of Skills. 1) Utilize human operation video to learn the skill sequence for task execution, decompose the video of the user’s progress so far into observations and human actions through segmentation, and input them along with task instruction into a pre-trained language model to predict the next step (patel2023pretrained). 2) LLM scores the skills in the skill library based on task instruction and the skills already executed, and the value function also scores the skills in the skill library based on observation images. The highest-scoring skill, obtained by multiplying the two scores, is selected as the next step (ahn2022can). The value function can consider multiple factors such as affordance, safety, user preference, and more (huang2023grounded), and these considerations can also be fine-tuning LLM (wu2023tidybot). 3) LLM assists the classical planner by translating task instruction into PDDL descriptions, sending them to the classical planner to generate a PDDL plan, and then translating the PDDL plan into a natural language plan using LLM (liu2023llm).
@@
As for the limitation of scene grounding, LLMs need to inquire about the scene representation to determine the availability, relationship and location of objects. NLMap (chen2023open) proposes an open-vocabulary, queryable semantic representation map built on ViLD and CLIP. This map outputs the pose of related objects based on task instruction, which are then handed over to the LLM for planning. Text2Motion (lin2023text2motion) incorporates a geometric value function on top of the value function, enabling the robot to select actions that adhere to geometric constraints based on scene descriptions. xu2023creative explore the possibility of teaching robots to creatively utilize tools within scenarios, which involve implicit physical limitations and require long-term planning. VILA (hu2023look) seamlessly incorporates perceptual data into ChatGPT-4V for its reasoning and planning processes, facilitating a deep comprehension of common sense knowledge within the visual domain, encompassing spatial arrangements and object characteristics. PHYSOBJECTS (gao2023physically) fine-tunes a VLM to enhance its understanding of physical object attributes, such as material. This integration of a physically informed VLM into an interactive framework with a LLM enhances task planning performance in tasks incorporating instruction related to physical object attributes. SpatialVLM (chen2024spatialvlm) and 3D-LLM (hong20233dllm) utilizes a 2D pre-trained VLM to train on collected 3D datasets, enhancing capabilities related to 3D tasks while maintaining the abilities of previous tasks.
@@
The hierarchy of skills possessed by LLMs or VLMs can be applied not only to single agent but also to multiple agents. SMART-LLM (kannan2023smart) utilizes LLM for the hierarchy of skills and allocates each task to every agent through the task assignment module.